OpenStudy (anonymous):

A large body collides with a stationary small body from the rear end. How do the bodies behave if the collision is inelastic?

OpenStudy (souvik):

some amount of kinetic energy transfer into sound or light or heat energy...

OpenStudy (anonymous):

Let us try to apply cons of momentum M=mass of large body m=mass of small body v=initial velocity of M V=final common velocity after inelastic coll Mv = (M+m)V V=Mv/(M+m) if M>>m V=Mv/M =v Which means that big body will still keep on cruising with its initial velocity dragging the little body with it.

OpenStudy (anonymous):

does it happen to move in an opposite direction?